# Totally odd immersions of complete graphs in graph products

### v1

@@ 2 Strong Product
Table 2: Rules for the definition of paths joining $(u_{i},v_{j^{\prime}})$ to $(u_{i^{\prime}},v_{j})$ in the case that $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$.
@@
We claim that due to the construction, the set of paths $\mathcal{M}$ formed by all the odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$, $M_{(i,j^{\prime})-(i^{\prime},j)}$ with $1\leq i\leq i^{\prime}\leq t$ and $1\leq j\leq j^{\prime}\leq s$ is pairwise edge-disjoint.
@@
Let us take the path $M_{(i,j)-(i,j^{\prime})}$ for some $1\leq i\leq t$ and $1\leq j<j^{\prime}\leq s$. Any edge $(x,y)-(x^{\prime},y^{\prime})$ from the path $M_{(i,j)-(i,j^{\prime})}$ satisfies that $x=x^{\prime}=u_{i}$ is a terminal, and $yy^{\prime}\in Q_{jj^{\prime}}$. So, by an exhaustive check we have that if $M_{(i,j)-(i,j^{\prime})}$ shares an edge with another path from $\mathcal{M}$, this path should be $M_{(i,q)-(i,q^{\prime})}$ for some $1\leq q<q^{\prime}\leq s$ with $\{q,q^{\prime}\}\neq\{j,j^{\prime}\}$. But then, $yy^{\prime}$ is an edge common in $Q_{jj^{\prime}}$ and $Q_{qq^{\prime}}$, which contradicts to the fact that $Q_{jj^{\prime}}$ and $Q_{qq^{\prime}}$ are paths from the totally odd strong immersion of $K_{s}$ in $H$. Analogously, we have the result for any path $M_{(i,j)-(i^{\prime},j)}$ for every $1\leq i<i^{\prime}\leq t$ and $1\leq j\leq s$.
@@
Now the only thing left is to show that $M_{(i_{1},j_{1})-(i_{1}^{\prime},j_{1}^{\prime})}$ and $M_{(i_{2},j_{2})-(i_{2}^{\prime},j_{2}^{\prime})}$ are edge disjoint, for any natural number $1\leq i_{1}\leq i_{1}^{\prime}\leq t,1\leq i_{2}<i_{2}^{\prime}\leq t,\leq j_{1},j_{1}^{\prime}(\neq j_{1})\leq s$ and $1\leq j_{2},j_{2}^{\prime}(\neq j_{2})\leq s$ By the sake of contradiction, we suppose $(x,y)-(x^{\prime},y^{\prime})$ is a common edge of the aforementioned paths. By an exhaustive check, we can ensure that $(x,y)-(x^{\prime},y^{\prime})$ satisfies that $xx^{\prime}\in E(P_{i_{1}i_{1}^{\prime}})$ and $yy^{\prime}\in E(Q_{j_{1}^{\prime}j_{1}})$ and at the same time $xx^{\prime}\in E(P_{i_{2}i_{2}^{\prime}})$ and $yy^{\prime}\in E(Q_{j_{2}^{\prime}j_{2}})$. However, $P_{i_{1}i_{1}^{\prime}}$ and $P_{i_{2}i_{2}^{\prime}}$ (resp., $Q_{j_{1}j_{1}^{\prime}}$ and $Q_{j_{2}j_{2}^{\prime}}$) are paths of the totally odd strong immersion of $K_{t}$ in $G$ (resp., of $K_{s}$ in $H$), and hence they are either edge-disjoint or equal. Thus, $\{i_{1},i_{1}^{\prime}\}=\{i_{2},i_{2}^{\prime}\}$ (thus $i_{1}=i_{2}$ and $i_{1}^{\prime}=i_{2}^{\prime}$) and $\{j_{1},j_{1}^{\prime}\}=\{j_{2},j_{2}^{\prime}\}$. In other words, $M_{(i_{2},j_{2})-(i_{2}^{\prime},j_{2}^{\prime})}$ is $M_{(i_{2},j_{2}^{\prime})-(i_{2}^{\prime},j_{2})}$. Again by an exhaustive check one can ensure that $M_{(i_{1},j_{1})-(i_{1}^{\prime},j_{1}^{\prime})}$ are $M_{(i_{2},j_{2}^{\prime})-(i_{2}^{\prime},j_{2})}$ edge disjoint paths. Hence we obtain a contradiction. ∎

### v4

@@ 2 Direct Product
Table 2: Rules for the definition of paths joining $(u_{i},v_{j^{\prime}})$ to $(u_{i^{\prime}},v_{j})$ in the case that $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$. In the subcase $k=0,\ell\geq 2$, the first coordinate alternates between $u_{i}$ and $u_{i^{\prime}}$, and the second coordinate moves along the path $Q_{jj^{\prime}}$ from $v_{j^{\prime}}$ to $v_{j}$. In the subcase $\ell=0,k\geq 2$, the second coordinate alternates between $v_{j^{\prime}}$ and $v_{j}$, and the first coordinate moves along the path $P_{ii^{\prime}}$.
Again, in this case, we can argue similarly that for each subcase, paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$ are vertex-disjoint. For the subcase $k=\ell$, in $M_{(i,j)-(i^{\prime},j^{\prime})}$ there are two types of vertices: the end vertices $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$ and vertices $(a_{p},b_{p})$. None of these vertices appear in $M_{(i,j^{\prime})-(i^{\prime},j)}$, since vertices in $M_{(i,j^{\prime})-(i^{\prime},j)}$ are of the following three types: the end vertices $(u_{i},v_{j^{\prime}})$ and $(u_{i^{\prime}},v_{j})$, and vertices $(a_{odd},b_{even})$, $(a_{even},b_{odd})$. For the subcase $k=0,\ell\geq 2$, in $M_{(i,j)-(i^{\prime},j^{\prime})}$ there are three types of vertices: the end vertices $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$, and vertices $(u_{i^{\prime}},b_{odd})$, $(u_{i},b_{even})$. None of these vertices appear in $M_{(i,j^{\prime})-(i^{\prime},j)}$, since vertices in $M_{(i,j^{\prime})-(i^{\prime},j)}$ are of the following three types: the end vertices $(u_{i},v_{j^{\prime}})$ and $(u_{i^{\prime}},v_{j})$, and vertices $(u_{i^{\prime}},b_{even})$, $(u_{i},b_{odd})$. Finally, for the subcase $\ell=0,k\geq 2$, in $M_{(i,j)-(i^{\prime},j^{\prime})}$ there are three types of vertices: the end vertices $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$, and vertices $(a_{odd},v_{j^{\prime}})$, $(a_{even},v_{j})$. None of these vertices appear in $M_{(i,j^{\prime})-(i^{\prime},j)}$, since vertices in $M_{(i,j^{\prime})-(i^{\prime},j)}$ are of the following three types: the end vertices $(u_{i},v_{j^{\prime}})$ and $(u_{i^{\prime}},v_{j})$, and vertices $(a_{odd},v_{j})$, $(a_{even},v_{j^{\prime}})$.
@@
We claim that the set of all odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$, $M_{(i,j^{\prime})-(i^{\prime},j)}$ with $1\leq i\leq i^{\prime}\leq t$ and $1\leq j\leq j^{\prime}\leq s$ is pairwise edge-disjoint.
@@
Let us take the path $M_{(i,j)-(i,j^{\prime})}$ for some $1\leq i\leq t$ and $1\leq j<j^{\prime}\leq s$. Note that every edge $(x,y)-(x^{\prime},y^{\prime})$ in the path $M_{(i,j)-(i,j^{\prime})}$ satisfies that $x=x^{\prime}=u_{i}$ and $yy^{\prime}\in Q_{jj^{\prime}}$. Thus, if $M_{(i,j)-(i,j^{\prime})}$ shares an edge with another path from $\mathcal{M}$, this path should be $M_{(i,q)-(i,q^{\prime})}$ for some $1\leq q<q^{\prime}\leq s$ with $\{q,q^{\prime}\}\neq\{j,j^{\prime}\}$. But then, $yy^{\prime}$ is an edge in common in $Q_{jj^{\prime}}$ and $Q_{qq^{\prime}}$, which contradicts the fact that $Q_{jj^{\prime}}$ and $Q_{qq^{\prime}}$ are paths from a totally odd strong immersion of $K_{s}$ in $H$. Analogously, the same holds for any path $M_{(i,j)-(i^{\prime},j)}$ for every $1\leq i<i^{\prime}\leq t$ and $1\leq j\leq s$.
It is left to show that the set of paths
$\{M_{(i,j)-(i^{\prime},j^{\prime})},M_{(i,j^{\prime})-(i^{\prime},j)}\mid 1\leq i<i^{\prime}\leq t,1\leq j<j^{\prime}\leq s\}$
@@
is pairwise edge disjoint. Let $M_{(i_{1},j_{1})-(i_{1}^{\prime},j_{1}^{\prime})}$ and $M_{(i_{2},j_{2})-(i_{2}^{\prime},j_{2}^{\prime})}$ distinct paths with $1\leq i_{1}<i_{1}^{\prime}\leq t,1\leq i_{2}<i_{2}^{\prime}\leq t,1\leq j_{1},j_{1}^{\prime}\leq s$ and $1\leq j_{2},j_{2}^{\prime}\leq s$, with $j_{1}\neq j_{1}^{\prime}$ and $j_{2}\neq j_{2}^{\prime}$. By the sake of contradiction, we suppose $(x,y)-(x^{\prime},y^{\prime})$ is a common edge of the aforementioned paths. Due to the construction of the paths we have that $xx^{\prime}\in E(P_{i_{1}i_{1}^{\prime}})$ and $yy^{\prime}\in E(Q_{j_{1}^{\prime}j_{1}})$ and at the same time $xx^{\prime}\in E(P_{i_{2}i_{2}^{\prime}})$ and $yy^{\prime}\in E(Q_{j_{2}^{\prime}j_{2}})$. However, $P_{i_{1}i_{1}^{\prime}}$ and $P_{i_{2}i_{2}^{\prime}}$ ($Q_{j_{1}j_{1}^{\prime}}$ and $Q_{j_{2}j_{2}^{\prime}}$, respectively) are paths of the totally odd strong immersion of $K_{t}$ in $G$ ($K_{s}$ in $H$, respectively), and hence, they are either edge-disjoint or equal. Thus, $P_{i_{1}i_{1}^{\prime}}=P_{i_{2}i_{2}^{\prime}}$ and $Q_{j_{1}j_{1}^{\prime}}=Q_{j_{2}j_{2}^{\prime}}$, and since $M_{(i_{1},j_{1})-(i_{1}^{\prime},j_{1}^{\prime})}$ and $M_{(i_{2},j_{2})-(i_{2}^{\prime},j_{2}^{\prime})}$ are distinct, then, $M_{(i_{2},j_{2})-(i_{2}^{\prime},j_{2}^{\prime})}$ is the path $M_{(i_{1},j_{1}^{\prime})-(i_{1}^{\prime},j_{1})}$. However, we have already proven that $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$ are edge-disjoint (actually, vertex-disjoint) for every $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$.
Now, we show that $\{\phi(c)\in\{(u_{i},v_{j})\mid 1\leq i\leq t,1\leq j\leq s\}\mid c\in\{c_{1},\ldots,c_{r}\}\}$ is the set of terminals of a totally odd strong immersion of $K_{r}$ in $G\times H$.
Suppose $\phi(c)=(u_{i},v_{j})$ and $\phi(c^{\prime})=(u_{i^{\prime}},v_{j^{\prime}})$ for some $1\leq i\leq i^{\prime}\leq t$ and $j,j^{\prime}\in\{1,\ldots,s\}$.
  • If $i\neq i^{\prime}$ and $j\neq j^{\prime}$, then for the path joining $\phi(c)$ and $\phi(c^{\prime})$ we consider $M_{(i,j)-(i^{\prime},j^{\prime})}$ or $M_{(i,j^{\prime})-(i^{\prime},j)}$ depending on whether $j<j^{\prime}$ or $j>j^{\prime}$.
  • If either $i=i^{\prime}$ or $j=j^{\prime}$, then first we take the odd path
    $(x_{i},y_{j})-(x_{i_{1}},y_{j_{1}})-(x_{i_{2}},y_{j_{2}})-\dots-(x_{i_{q}},y_{j_{q}})-(x_{i^{\prime}},y_{j^{\prime}})$
    joining $(x_{i},y_{j})$ to $(x_{i^{\prime}},y_{j^{\prime}})$ in the totally odd strong immersion of $K_{r}$ in $K_{t}\times K_{s}$. Note that each edge $(x,y)-(x^{\prime},y^{\prime})$ from this path satisfies that $x\neq x^{\prime}$ and $y\neq y^{\prime}$ due to the definition of direct product. Recall that $\phi(x_{i},y_{j})=(u_{i},v_{j})$ for $1\leq i\leq t$ and $1\leq j\leq s$. See Figure 1 for an illustration of this case. Thus, we can now define the odd path of the immersion of $K_{r}$ in $G\times H$ joining $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ as the union of the following odd paths
    $M^{*}_{(i,j)-({i_{1}},{j_{1}})}-M^{*}_{(i_{1},j_{1})-(i_{2},j_{2})}-M^{*}_{({i_{2}},{j_{2}})-({i_{3}},{j_{3}})}-\dots-M^{*}_{(i_{q-1},j_{q-1})-(i_{q},j_{q})}-M^{*}_{(i_{q},j_{q})-(i^{\prime},j^{\prime})}$
    where
[…]
Note that these odd paths are edge-disjoint as
$\{M_{(i,j)-(i^{\prime},j^{\prime})},M_{(i,j^{\prime})-(i^{\prime},j)}\mid 1\leq i<i^{\prime}\leq t,1\leq j<j^{\prime}\leq s\}$
is a set of pairwise edge-disjoint odd paths in $G\times H$, as proven in the beginning of the proof. Therefore, $G\times H$ has $K_{r}$ as a totally odd strong immersion.
Figure 1: Let us suppose that the odd path of the totally odd strong immersion of $K_{r}$ in $K_{t}\times K_{s}$ that joins $(x_{i},y_{j})$ and $(x_{i^{\prime}},y_{j})$ is given by the sequence of vertices $(x_{i},y_{j})-(x_{i_{1}},y_{j_{1}})-(x_{i_{2}},y_{j_{2}})-(x_{i^{\prime}},y_{j})$. The illustration shows the result of the union of the paths $M^{*}_{(i,j)-({i_{1}},{j_{1}})}$ in green, $M^{*}_{({i_{1}},{j_{1}})-({i_{2}},{j_{2}})}$ in red and $M^{*}_{({i_{2}},{j_{2}})-(i^{\prime},j)}$ in dark red, which corresponds to the odd path joining $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j})$ in the totally odd strong immersion of $K_{r}$ in $G\times H$.
Corollary 2.2.
The direct product of two graphs is not a minimum counterexample to Conjecture 1.3.
Proof.
Let us assume that there are graphs $G$ and $H$ so that $G\times H$ is a minimum counterexample to Conjecture 1.3, and that $\mathrm{toi}(G)=t,\mathrm{toi}(H)=s$. Since $\times$ is commutative, we can assume that $t\leq s$. Due to Theorem 2.1, we have that $\mathrm{toi}(G\times H)\geq\mathrm{toi}(K_{t}\times K_{s})$. Note that If $V(K_{t})=\{x_{1},x_{2},x_{3},\dots,x_{t}\}$, and $V(K_{s})=\{y_{1},y_{2},y_{3},\dots,y_{s}\}$, then the set $\{(x_{i},y_{i})\in V(K_{t}\times K_{s}):i\in\{1,\dots,t\}\}$ induces $K_{t}$ as a subgraph in $K_{t}\times K_{s}$, and hence $\mathrm{toi}(K_{t}\times K_{s})\geq t=\min\{\mathrm{toi}(G),\mathrm{toi}(H)\}$. Since $G$ and $H$ are not counterexamples to Conjecture 1.3, we have $\mathrm{toi}(G)\geq\chi(G)$, $\mathrm{toi}(H)\geq\chi(H)$, and hence $\min\{\mathrm{toi}(G),\mathrm{toi}(H)\}\geq\min\{\chi(G),\chi(H)\}$. Furthermore, we have that $\min\{\chi(G),\chi(H)\}\geq\chi(G\times H)$, which finally yields a contradiction. ∎
Our next result regarding the paramenter $\mathrm{toi}$ for the direct product of graphs tells us that the largest totally odd strong immersion in $K_{2t}\times K_{s}$ of a complete graph has at least $ts$ vertices. We think, it is possible to improve the bound of Theorem 2.3, however many more ad-hock paths would be needed. In [5], the authors obtain (in Theorem 18) the largest strong immersion of the product of two complete graphs, where most of the paths are of even length. Here, we construct paths of length one and three.
Theorem 2.3.
$\mathrm{toi}(K_{2t}\times K_{s})\geq ts$, for all natural numbers $t\geq 6$ and $s\geq 5$.
Proof.
Let $\{u_{1},u_{2},u_{3},\dots u_{2t}\}$ and $\{v_{1},v_{2},v_{3}\dots v_{s}\}$ be the set of vertices in $K_{2t}$ and $K_{s}$, respectively. We claim that $K_{2t}\times K_{s}$ has a $K_{ts}$ as a totally odd strong immersion with the set of terminals $A:=\{(u_{2i-1},v_{j})\mid 1\leq i\leq t,1\leq j\leq s\}$. First, note that two elements of $A$ that differ in both coordinates are adjacent in the product graph. Let us denote this set of edges by $\mathcal{E}_{s}$. Thus, we only need to connect pairs of vertices that differ in exactly one coordinate. This connection is made through edge-disjoint paths of length three, and we use edges in $E(K_{2t}\times K_{s})\setminus\mathcal{E}_{s}$. In order to prove that the paths that we define are pairwise edge-disjoint, we partition $E(K_{2t}\times K_{s})\setminus\mathcal{E}_{s}$ into classes according to the function $f:E(K_{2t}\times K_{s})\setminus\mathcal{E}_{s}\longrightarrow\mathbb{Z}^{3}$ defined as
$f(u_{i},v_{j})(u_{i^{\prime}},v_{j^{\prime}})=(i~\mod(2),i-i^{\prime},j^{\prime}-j)\qquad\text{whenever}\qquad j<j^{\prime}.$
An example is illustrated in Figure 2.
[…]
Figure 2: The edge $e_{1}$ in the left hand side belongs to the class (1,1,5)(1,-1,5) and the edge $e_{2}$ in the right hand side belongs to the class $(0,2,7)$.
Since $f$ is well-defined, we have that
Claim 2.4.
If $f(e)\neq f(e^{\prime})$, then $e\neq e^{\prime}$. In other words, edges in different classes are distinct.
Furthermore, each non-empty class defined according to the function $f$ contains all translations of an edge. In this context, let $e=(u_{i},v_{j})(u_{i^{\prime}},v_{j^{\prime}})$ and $e^{\prime}=(u_{k},v_{l})(u_{k^{\prime}},v_{l^{\prime}})$ be edges, we say that $e^{\prime}$ is a translation of $e$ if there exist two integers $a,b\geq 0$ such that $k=i+2a$, $k^{\prime}=i^{\prime}+2a$, $l=j+b$ and $l^{\prime}=j^{\prime}+b$.
Claim 2.5.
Let $e,e^{\prime}$ be two edges in $E(K_{2t}\times K_{s})\setminus\mathcal{E}_{t}$. We have that $f(e)=f(e^{\prime})$ if only if either $e=e^{\prime}$ or $e^{\prime}$ is a translation of $e$.
Proof.
The only if direction is trivial. Suppose that $f(e)=f(e^{\prime})$, if $e=e^{\prime}$ we are ready. Suppose that $e\neq e^{\prime}$, we have to prove that $e^{\prime}=(u_{k},v_{l})(u_{k^{\prime}},v_{l^{\prime}})$ is a translation of $e=(u_{i},v_{j})(u_{i^{\prime}},v_{j^{\prime}})$. First, $i~\mod(2)=k~\mod(2)$, it means that there is a $x\geq 0$ such that $i=k+2x$. Now, we have $i^{\prime}-i=k^{\prime}-k$, taking $\mod(2)$ on both sides, we get $i^{\prime}~\mod(2)=k^{\prime}~\mod(2)$, then there is a $y$ such that $i^{\prime}=k^{\prime}+2y$, it is not hard to see that $x=y$. Suppose that there is no integer $b\geq 0$ such that $j^{\prime}-j\neq b$ and $l^{\prime}-l\neq b$, then $j^{\prime}-j\neq l^{\prime}-l$, contradiction. Finally, $e^{\prime}$ is a translation of $e$. ∎
In the following, we describe the paths for the totally odd strong immersion. We divide the description into cases. In each case, we define a general "pattern/shape" for the paths (given by the classes of their edges) and take translations of them to cover all pairs of vertices within the case. When in the translation, we move the edges of the paths "too much to the right/left" or "too much down/up", we end up having edges which turn around the grid and as a consequence their classes change; that is the main reason for having sub-cases in each case. However, it is convenient for the reader to keep in mind that the "pattern/shape" of the paths is kept. Furthermore, in each case, we make sure that the edges of the paths described in different cases belong to different classes, and hence the paths described in different cases are naturally edge-disjoint.
  • Case 1
    For $1\leq i\leq t-1$ and $1\leq j<j^{\prime}\leq s$, we define the path joining (2i1,j)(2i-1,j) and $(2i-1,j^{\prime})$ as
    $P_{(2i-1,j)-(2i-1,j^{\prime})}:=(u_{2i-1},v_{j})-(u_{2i},v_{j^{\prime}})-(u_{2i+2},v_{j})-(u_{2i-1},v_{j^{\prime}})$
    We find that the first edge of the path $P_{(2i-1,j)-(2i-1,j^{\prime})}$ is $(u_{2i-1},v_{j})-(u_{2i},v_{j^{\prime}})$ and hence it is in the class $(1,-1,j^{\prime}-j)$, the second edge $(u_{2i},v_{j^{\prime}})-(u_{2i+2},v_{j})$ is in the class $(0,2,j^{\prime}-j)$ and the third edge $(u_{2i+2},v_{j})-(u_{2i-1},v_{j^{\prime}})$ is in the class $(0,3,j^{\prime}-j)$, where $1\leq j^{\prime}-j\leq s-1$.
    In the case of $i=t$ and $1\leq j<j^{\prime}\leq s$, the path joining (2t1,j)(2t-1,j) and $(2t-1,j^{\prime})$ is defined as
    $P_{(2t-1,j)-(2t-1,j^{\prime})}:=(u_{2t-1},v_{j})-(u_{2t},v_{j^{\prime}})-(u_{2},v_{j})-(u_{2t-1},v_{j^{\prime}})$
    and the first edge is in the class $(1,-1,j^{\prime}-j)$, the second edge in the class $(0,2-2t,j^{\prime}-j)$ and the third edge in the class $(0,3-2t,j^{\prime}-j)$, where $1\leq j^{\prime}-j\leq s-1$.
    Let $\mathcal{P}$ denote the class of all described paths. Now, if we consider two edges $e$, $e^{\prime}$ in different paths $P,P^{\prime}\in\mathcal{P}$ that belong to the same class, then they are both the first, second or third edges of the respective path. As the paths are different, they are joining different pairs of vertices, and hence the edges are joining different pairs of vertices as well. For a better visualization of the odd paths, see Figure 6.
    Figure 3: When $j=1$ and $j^{\prime}\in\{2,3,4,5\}$
    Figure 4: When $j=2$ and $j^{\prime}\in\{3,4,5\}$
    Figure 5: When $j=3$ and $j^{\prime}\in\{4,5\}$
    Figure 6: The picture represents the odd paths in the case 1, when $i\in\{1,t-1\}$, $j,j^{\prime}\in\{1,2,3,4,5\}$ with $j<j^{\prime}$. The terminals of the totally odd strong immersion of the $K_{ts}$ are in red while the others are in black.
  • Case 2
    For $1\leq i\leq t-1$ and $1\leq j\leq s-2$, the paths are defined as follows.
    $Q_{(2i-1,j)-(2i+1,j)}=(u_{2i-1},v_{j})-(u_{2i+2},v_{j+2})-(u_{2i},v_{j+1})-(u_{2i+1},v_{j})$
[…]
    For $1\leq i\leq t-3$ and $s-1\leq j\leq s$, the paths are defined as follows.
    $Q_{(2i-1,j)-(2i+1,j)}=(u_{2i-1},v_{j})-(u_{2i},v_{j-1})-(u_{2i+6},v_{j-(s-2)})-(u_{2i+1},v_{j})$
[…]
    For $t-2\leq i\leq t-1$ and $j\in\{s-1,s\}$, the paths are defined as follows.
    $Q_{(2i-1,j)-(2i+1,j)}=(u_{2i-1},v_{j})-(u_{2i-4},v_{j-(s-2)})-(u_{2i},v_{j-(s-3)})-(u_{2i+1},v_{j})$
[…]
    Again, from the definition it is clear that edges from different paths in the same class are translations of each other (and hence not equal), since they join different pairs of vertices. For an example of some odd path, we refer to the Figure 7.
    Figure 7: The picture shows the first odd paths described in case 2, when $i\in\{1,2,\dots,t-1\}$ and $j\in\{1,3\}$. For instance, the first blue odd path is $Q_{(1,1)-(1,3)}=(u_{1},v_{1})-(u_{4},v_{3})-(u_{2},v_{2})-(u_{3},v_{1})$.
  • Case 3
    For $1\leq i<i^{\prime}\leq t$ with $i^{\prime}\neq i+1$ and for $1\leq j\leq s-2$. The paths are defined as follows.
    $Q_{(2i-1,j)-(2i^{\prime}-1,j)}=(u_{2i-1},v_{j})-(u_{2i^{\prime}},v_{j+1})-(u_{2i},v_{j+2})-(u_{2i^{\prime}-1},v_{j})$
[…]
    For $1\leq i<i^{\prime}\leq t$ with $i^{\prime}\neq i+1$ and for j=s1j=s-1, , except for when $i=1$ and $i^{\prime}=t$ simultaneously. The paths are as follows.
    $Q_{(2i-1,s-1)-(2i^{\prime}-1,s-1)}=(u_{2i-1},v_{s-1})-(u_{2i^{\prime}},v_{s})-(u_{2i},v_{1})-(u_{2i^{\prime}-1},v_{s-1})$
[…]
    For $1\leq i<i^{\prime}\leq t$ with $i^{\prime}\neq i+1$ and for $j=s$, except for when $i=1$ and $i^{\prime}=t$ simultaneously. The path is defined as follows.
    $Q_{(2i-1,s)-(2i^{\prime}-1,s)}=(u_{2i-1},v_{s})-(u_{2i^{\prime}},v_{1})-(u_{2i},v_{2})-(u_{2i^{\prime}-1},v_{s})$
[…]
    Finally, for $j\in\{s-1,s\}$ and when $i=1$ and $i^{\prime}=t$ simultaneously. The paths are defined as follows.
    $Q_{(1,j)-(2t-1,j)}=(u_{1},v_{j})-(u_{2t},v_{j-1})-(u_{2t-6},v_{j-2})-(u_{2t-1},v_{j})$
[…]